# DemoTuner: Efficient DBMS Knobs Tuning via LLM-Assisted Demonstration Reinforcement Learning

### v1

@@ Abstract
The performance of modern DBMSs such as MySQL and PostgreSQL heavily depends on the configuration of performance-critical knobs. Manual tuning these knobs is laborious and inefficient due to the complex and high-dimensional nature of the configuration space. Among the automated tuning methods, reinforcement learning (RL)-based methods have recently sought to improve the DBMS knobs tuning process from several different perspectives. However, they still encounter challenges with slow convergence speed during offline training. In this paper, we mainly focus on how to leverage the valuable tuning hints contained in various textual documents such as DBMS manuals and web forums to improve the offline training of RL-based methods. To this end, we propose an efficient DBMS knobs tuning framework named DemoTuner via a novel LLM-assisted demonstration reinforcement learning method. Specifically, to comprehensively and accurately mine tuning hints from documents, we design a structured chain of thought prompt to employ LLMs to conduct a condition-aware tuning hints extraction task. To effectively integrate the mined tuning hints into RL agent training, we propose a hint-aware demonstration reinforcement learning algorithm HA-DDPGfD in DemoTuner. As far as we know, DemoTuner is the first work to introduce the demonstration reinforcement learning algorithm for DBMS knobs tuning. Experimental evaluations conducted on MySQL and PostgreSQL across various workloads demonstrate the significant advantages of DemoTuner in both performance improvement and online tuning cost reduction over three representative baselines including DB-BERT, GPTuner and CDBTune. Additionally, DemoTuner also exhibits superior adaptability to application scenarios with unknown workloads.
@@ I Introduction
Modern database management systems (DBMSs) such as MySQL and PostgreSQL provide many performance-critical knobs and their performance (throughput, latency, etc.) heavily relies on the configuration of these knobs. Considering the complex relationship between performance and knobs as well as the high dimensionality of configuration space, manually tuning the performance-critical knobs by database administrators (DBAs) is both labor-intensive and time-consuming. Consequently, how to automatically tune these knobs to improve the performance of DBMSs has attracted considerable interest from both academic and industrial sectors [duan2009tuning, van2017ottertune, cereda2021cgptuner, zhang2022towards, zhang2021restune, zhang2019CDBTune, xiong2017ath, bao2018autoconfig, li2019qtune, ge2021watuning, yang2024vdtuner] in the past few years.
@@
Broadly speaking, state-of-the-art automated knobs tuning methods can be categorized into search-based, model-based and reinforcement learning (RL)-based. Specifically, search-based methods [duan2009tuning, cereda2021cgptuner, zhang2022towards, zhang2021restune, yang2024vdtuner] employ techniques such as Bayesian optimization [shahriari2015taking, hutter2011sequential, cowen2020hebo] to iteratively explore the configuration space to identify optimal knob settings in an online manner. However, since these methods do not effectively utilize historical tuning experiences and lack awareness of the runtime environment, they usually require extensive time-consuming configuration evaluations and must restart from scratch when workload or hardware changes happen. Model-based methods [van2017ottertune, xiong2017ath, bao2018autoconfig] initially train an offline model to predict the relationship between knob configurations and performance, and then conduct online knobs tuning based solely on this predictive model. Consequently, these approaches usually require a large number of high-quality samples to develop an accurate performance prediction model. Moreover, the poor adaptability of conventional machine learning algorithms often necessitates model retraining when changes in workloads or hardware resources occur. In contrast, RL-based methods [zhang2019CDBTune, li2019qtune, ge2021watuning] learn policies that map current runtime states to an action (i.e., configuration of knobs) in a trial-and-error manner during the offline training phase. After training, the agent only needs a few sequential online tuning iterations to adapt to the practical application scenario so as to recommend satisfying knob configurations.
@@
Recent RL-based methods have sought to improve the DBMS knobs tuning process from several perspectives [cai2022hunter, dou2024deepcat+, henderson2022blutune]. However, they still encounter challenges with slow convergence speed during offline training, which negatively impacts both the effectiveness (performance improvement) and efficiency (tuning costs) of the subsequent online tuning phase. On the one hand, without domain knowledge for knobs tuning under current DBMS runtime environment, the poor initial policy of a RL agent heavily impedes its offline training process. On the other hand, high-performing knob configurations are very sparse and distributed in the entire configuration space, potentially causing the RL agent to experience a prolonged period without significant positive feedback. Fortunately, current DBMS manuals, web forums and other documents offer a wealth of information about knobs tuning [trummer2023db][lao2025gptuner]. Either derived from domain knowledge or validated from practice, these valuable tuning hints can be utilized to enhance the offline training phase of RL-based knobs tuning methods, thereby improving the performance of their online tuning phase. Nevertheless, there are still two major pending challenges to achieve this promising objective:
@@
CH1: How to comprehensively and accurately mine runtime-aware tuning hints from various documents? Previous natural language processing (NLP)-enhanced DBMS knobs tuning methods such as DB-BERT [trummer2023db] only focus on textual information regarding the selection and configuration of performance-critical knobs contained in the related documents. However, both the importance of knobs and their optimal configuration are highly dependent on the runtime environment such as workload characteristics, system states and resource usage, directly ignoring these runtime information may lead to the inappropriate usage of tuning hints. For instance, as shown in Table I, text about the MySQL knob innodb_log_file_size recommends setting it to 4 GB for a write-heavy workload rather than the common 512 MB. Additionally, some texts do not recommend any explicit knob values, instead, they only provide implicit tuning advice. For example, although the text about sort_buffer_size does not include specific value, however, we can still infer that the knob value should be increased when workload is read-heavy or involves many sort operations. This direction guidance can help prevent the RL agent from taking incorrect actions. Regrettably, such valuable implicit tuning hints have been completely overlooked by previous studies [trummer2023db, lao2025gptuner].
@@
  Text
Starting with innodb_log_file_size
= 512M (giving 1GB of redo logs
) should give you plenty of room
for writes. If you know your appli-
cation is write-intensive, you can
start with innodb_log_file_size =
4G.
sort_buffer_size: For a read-
heavy workload (again in
which tables don’t have
indexes), sort queries require
memory to complete
a full sort.
@@
Env   / write heavy
read heavy
no index
sort operation
CH2: How to effectively integrate the mined tuning hints into the RL agent training phase? The standard RL training process involves an agent interacting with the environment to collect transitions through trial-and-error to determine optimal actions based on states, but it lacks support for the integration of domain knowledge. To our knowledge, DB-BERT [trummer2023db] is currently the only attempt to merge tuning hints with RL for DBMS knobs tuning. This method utilizes a RL agent to translate tuning hints into knob values and prioritize hints according to runtime environmental feedback. However, in addition to the above mentioned shortcomings in tuning hints extraction, the knob configurations in DB-BERT are generated totally driven by tuning hints with a very limited range. On the one hand, the benefits of some tuning hints tend to diminish along with knobs tuning iterations and may even deviate from the intended tuning target. On the other hand, transitions with real environmental feedback may provide more information than the static domain knowledge stored in tuning hints. As shown in Figure 1, compared with the pure DDPG-based method CDBTune [zhang2019CDBTune], DB-BERT [trummer2023db] has an obviously larger convergence rate but gradually loses its competitiveness with ongoing training steps. Therefore, it is crucial to effectively incorporate the extracted tuning hints throughout the entire RL training phase to improve both the initial and long-term training performance of the tuning agent.
Figure 1: Current best performance achieved by DB-BERT[trummer2023db] and CDBTune[zhang2019CDBTune] along with their RL agents training steps.
@@
  • We conduct a series of experiments on different DBMSs with various workloads. Experimental results successfully verify the advantages of DemoTuner over the baseline methods. The source code of DemoTuner is available at [DemoTuner2025Codes].
@@ II-A RL-based DBMS Knobs Tuning
Knobs in DBMSs are closely related to their memory allocation, disk I/O, concurrency control and other factors. Knobs tuning means adjusting the values of knobs to improve the performance under specific execution environment. Reinforcement learning [sutton2018reinforcement] is a type of machine learning algorithms where an agent learns a policy to make decisions by performing actions in an environment to maximize the cumulative reward. In the context of DBMS knobs tuning, RL-based methods [zhang2019CDBTune, li2019qtune, ge2021watuning, li2023auto] treat it as a trial-and-error process based on system feedback and exhibit a degree of adaptability to unknown runtime environments. Although recent studies [cai2022hunter, dou2024deepcat+, henderson2022blutune, mai2024ctuner, wang2021udo, zhang2023unified] have enhanced the RL-based methods from various angles, training reinforcement learning models is still significant costly, especially in the field of knobs tuning, where each configuration evaluation for transition collection may require significant time consumption. The offline agent training usually suffers from the poor initial policy since the high-performing configurations are very sparse in the configuration space.
@@ II-B Tuning Hints Extraction
Manuals, web forums and other texts often offer a wealth of domain knowledge about knob settings for DBMSs. DB-BERT [trummer2023db] leverages BERT to mine these tuning hints in the form of <<knob, rec_value>>, where rec_value is the recommended value for the corresponding knobknob. However, as shown in Table I, since some texts only provide implicit tuning advice such as increasing and decreasing the knob, the extracted tuning hints should contain these tuning direction information. Besides, we should also pay attention to their application condition if we want accurately using the extracted tuning hints. Towards this target, we extend the definition of tuning hint as a recommended adjustment action for a knob under a specific runtime environment (i.e., application condition). Specifically, a condition-aware tuning hint can be described as a triple <<knob, rec_action, condition_vector>>, where condition_vector is a vector describes the applicable runtime environment of rec_action. We will give a detailed description about the definition of condition_vector and how to extract condition-aware tuning hints in Section IV.
@@ II-C LLM-Assisted and LLM-Driven Knobs Tuning
As the pioneering work of NLP-enhanced DBMS tuning methods, DB-BERT [trummer2023db] first utilizes BERT to extract the tuning hints from texts, and then leverages RL to deal with these tuning hints to generate knob values in an iterative manner. Although DB-BERT can incorporate domain knowledge into the knobs tuning task, however, it totally neglects the application condition of tuning hints, which may mislead the tuning agent. In addition, it is only able to explore a limited range around the recommended knob values by tuning hints, which may miss the opportunities to further improve the performance. Finally, as shown in Figure 1, totally driven by tuning hints may lose competitiveness along with training steps since the tuning hints’ utilities in performance improvement are not static. Considering the advantages of large language models (LLMs) in semantic understanding, reasoning and domain knowledge, several recent works begin to use LLMs for DBMS knobs tuning. We broadly categorize them into LLM-assisted [lao2025gptuner, huang2024llmtune] and LLM-driven [fan2024latuner, li2024large, thakkar2025elmo, giannakouris2025lambda]. The LLM-assisted methods usually employ a LLM to extract or output tuning hints and then conduct a traditional knobs tuning procedure. For instance, GPTuner [lao2025gptuner] uses GPT-4 to select performance-critical knobs and extract the suggested value ranges, and then conduct a Bayesian optimization (BO)-based knobs tuning process. However, tuning hints extracted by the LLMs are not always reliable [liang2025same], relying solely on GPT’s judgments does not ensure the alignment and accuracy of domain knowledge with the current environment since BO inherently lacks the ability to interact with runtime environment. We also employ LLMs to assist the tuning hints extraction task in DemoTuner. On the other hand, LLM-driven methods design prompts for the whole knobs tuning workflow including knobs selection, knobs range determination and knobs configuration. While promising, relying solely on LLMs to conduct knobs tuning is still full of challenges at the present stage considering the hallucination and domain adaptation issues.
@@ IV-B LLM-based Tuning Hints Extraction
To effectively mine tuning hints described as <<knob, rec_action, condition_vector>> from the collected source texts, we choose to utilize GPT-3.5-Turbo [gpt2025] after evaluating a range of advanced large language models (LLMs) considering the tradeoff between costs and practical effectiveness. How to select the best LLM is out of the scope of this paper. Figure 3 illustrates the prompt structure, which includes four critical elements: the model’s role and expertise, the input source text, the output format, and a task description following the chain of thought (CoT) framework [wei2022chain]. Specifically, the CoT process can be divided into three steps. The first involves the extraction of knobs specific to the target DBMSs such as MySQL or PostgreSQL. The second focuses on extracting recommended values of knobs. In cases where no specific value is provided, the model is required to infer whether the knob’s value should be increased or decreased. The third targets to extract the corresponding circumstances, which can be utilized to identify the application condition vector of this tuning hint. Incorporating the phrase “Let’s think step by step” to split up the task helps guide GPT’s inference process. Besides, placing constraints directly within each step improves GPT’s comprehension and execution by keeping the task and constraints closely aligned. These prompts sharpen GPT’s focus on DBMS knobs tuning and structure the task into clear steps, thus enhancing the extraction accuracy [kojima2022LLMZeroShot]. Notably, GPT’s responses sometimes diverge from the prescribed format and exhibit variability. In such cases, we employ GPT to further analyze the deviant text until it aligns with our specified processing standards.
@@ V-B Discussions About Demonstration Reinforcement Learning
During the training process of DRL, the policy network learns to select actions by optimizing its parameters based on environmental feedback, while the value function estimates the expected rewards for these actions, guiding the policy network through transitions to maximize cumulative rewards. Without domain knowledge, the training process usually has to start from scratch and thus suffers from the cold start problem. Unfortunately, conventional DRL algorithms including Deep Q-Networks (DQN) [mnih2015human], Policy Gradients [schulman2015high] and Actor-Critic [lillicrap2015continuous] cannot directly integrate domain knowledge into the training phase. To address this issue, demonstration reinforcement learning algorithms such as DDPGfD (Deep Deterministic Policy Gradient from Demonstrations) have been recently developed [vecerik2017leveraging]. These algorithms can integrate demonstration experiences into the learning process, effectively bridging the gap between pure reinforcement learning and imitation learning.

### v2

# DemoTuner: Automatic Performance Tuning for Database Management Systems Based on Demonstration Reinforcement Learning

@@ Abstract
The performance of modern DBMSs such as MySQL and PostgreSQL heavily depends on the configuration of performance-critical knobs. Manual tuning these knobs is laborious and inefficient due to the complex and high-dimensional nature of the configuration space. Among the automated tuning methods, reinforcement learning (RL)-based methods have recently sought to improve the DBMS knobs tuning process from several different perspectives. However, they still encounter challenges with slow convergence speed during offline training. In this paper, we mainly focus on how to leverage the valuable tuning hints contained in various textual documents such as DBMS manuals and web forums to improve the offline training of RL-based methods. To this end, we propose an efficient DBMS knobs tuning framework named DemoTuner via a novel LLM-assisted demonstration reinforcement learning method. Specifically, to comprehensively and accurately mine tuning hints from documents, we design a structured chain of thought prompt to employ LLMs to conduct a condition-aware tuning hints extraction task. To effectively integrate the mined tuning hints into RL agent training, we propose a hint-aware demonstration reinforcement learning algorithm HA-DDPGfD in DemoTuner. As far as we know, DemoTuner is the first work to introduce the demonstration reinforcement learning algorithm for DBMS knobs tuning. Experimental evaluations conducted on MySQL and PostgreSQL across various workloads demonstrate that DemoTuner achieves performance gains of up to 44.01% for MySQL and 39.95% for PostgreSQL over default configurations. Compared with three representative baseline methods including DB-BERT, GPTuner and CDBTune, DemoTuner is able to further reduce the execution time by up to 10.03%, while always consuming the least online tuning cost. Additionally, DemoTuner also exhibits superior adaptability to application scenarios with unknown workloads.
@@ I Introduction
Modern database management systems (DBMSs) such as MySQL and PostgreSQL provide many performance-critical knobs and their performance (throughput, latency, etc.) heavily relies on the configuration of these knobs. Considering the complex relationship between performance and knobs as well as the high dimensionality of configuration space, manually tuning the performance-critical knobs by database administrators (DBAs) is both labor-intensive and time-consuming. Consequently, how to automatically tune these knobs to improve the performance of DBMSs has attracted considerable interest from both academic and industrial sectors [10, 34, 3, 44, 43, 41, 38, 1, 19, 12, 40] in the past few years.
@@
Broadly speaking, state-of-the-art automated knobs tuning methods can be categorized into search-based, model-based and reinforcement learning (RL)-based. Specifically, search-based methods [10, 3, 44, 43, 40] employ techniques such as Bayesian optimization [30, 16, 6] to iteratively explore the configuration space to identify optimal knob settings in an online manner. However, since these methods do not effectively utilize historical tuning experiences and lack awareness of the runtime environment, they usually require extensive time-consuming configuration evaluations and must restart from scratch when workload or hardware changes happen. Model-based methods [34, 38, 1] initially train an offline model to predict the relationship between knob configurations and performance, and then conduct online knobs tuning based solely on this predictive model. Consequently, these approaches usually require a large number of high-quality samples to develop an accurate performance prediction model. Moreover, the poor adaptability of conventional machine learning algorithms often necessitates model retraining when changes in workloads or hardware resources occur. In contrast, RL-based methods [41, 19, 12] learn policies that map current runtime states to an action (i.e., configuration of knobs) in a trial-and-error manner during the offline training phase. After training, the agent only needs a few sequential online tuning iterations to adapt to the practical application scenario so as to recommend satisfying knob configurations.
@@
Recent RL-based methods have sought to improve the DBMS knobs tuning process from several perspectives [2, 9, 14]. However, they still encounter challenges with slow convergence speed during offline training, which negatively impacts both the effectiveness (performance improvement) and efficiency (tuning costs) of the subsequent online tuning phase. On the one hand, without domain knowledge for knobs tuning under current DBMS runtime environment, the poor initial policy of a RL agent heavily impedes its offline training process. On the other hand, high-performing knob configurations are very sparse and distributed in the entire configuration space, potentially causing the RL agent to experience a prolonged period without significant positive feedback. Fortunately, current DBMS manuals, web forums and other documents offer a wealth of information about knobs tuning [33][18]. Either derived from domain knowledge or validated from practice, these valuable tuning hints can be utilized to enhance the offline training phase of RL-based knobs tuning methods, thereby improving the performance of their online tuning phase. Nevertheless, there are still two major pending challenges to achieve this promising objective:
@@
CH1: How to comprehensively and accurately mine runtime-aware tuning hints from various documents? Previous natural language processing (NLP)-enhanced DBMS knobs tuning methods such as DB-BERT [33] only focus on textual information regarding the selection and configuration of performance-critical knobs contained in the related documents. However, both the importance of knobs and their optimal configuration are highly dependent on the runtime environment such as workload characteristics, system states and resource usage, directly ignoring these runtime information may lead to the inappropriate usage of tuning hints. For instance, as shown in Table I, text about the MySQL knob innodb_log_file_size recommends setting it to 4 GB for a write-heavy workload rather than the common 512 MB. Additionally, some texts do not recommend any explicit knob values, instead, they only provide implicit tuning advice. For example, although the text about sort_buffer_size does not include specific value, however, we can still infer that the knob value should be increased when workload is read-heavy or involves many sort operations. This direction guidance can help prevent the RL agent from taking incorrect actions. Regrettably, such valuable implicit tuning hints have been completely overlooked by previous studies [33, 18].
@@
  Text   Starting with innodb_log_file_size = 512M (giving 1GB of redo logs ) should give you plenty of room for writes. If you know your appli- cation is write-intensive, you can start with innodb_log_file_size = 4G.   sort_buffer_size: For a read- heavy workload (again in which tables don’t have indexes), sort queries require memory to complete a full sort.
@@
Env   / write heavy   read heavy no index sort operation
CH2: How to effectively integrate the mined tuning hints into the RL agent training phase? The standard RL training process involves an agent interacting with the environment to collect transitions through trial-and-error to determine optimal actions based on states, but it lacks support for the integration of domain knowledge. To our knowledge, DB-BERT [33] is currently the only attempt to merge tuning hints with RL for DBMS knobs tuning. This method utilizes a RL agent to translate tuning hints into knob values and prioritize hints according to runtime environmental feedback. However, in addition to the above mentioned shortcomings in tuning hints extraction, the knob configurations in DB-BERT are generated totally driven by tuning hints with a very limited range. On the one hand, the benefits of some tuning hints tend to diminish along with knobs tuning iterations and may even deviate from the intended tuning target. On the other hand, transitions with real environmental feedback may provide more information than the static domain knowledge stored in tuning hints. As shown in Figure 1, compared with the pure DDPG-based method CDBTune [41], DB-BERT [33] has an obviously larger convergence rate but gradually loses its competitiveness with ongoing training steps. Therefore, it is crucial to effectively incorporate the extracted tuning hints throughout the entire RL training phase to improve both the initial and long-term training performance of the tuning agent.
Figure 1: Current best performance achieved by DB-BERT[33] and CDBTune[41] along with their RL agents training steps.
@@
  • We conduct a series of experiments on different DBMSs with various workloads. Experimental results successfully verify the advantages of DemoTuner over the baseline methods. The source code of DemoTuner is available at [7].
@@ II-A RL-based DBMS Knobs Tuning
Knobs in DBMSs are closely related to their memory allocation, disk I/O, concurrency control and other factors. Knobs tuning means adjusting the values of knobs to improve the performance under specific execution environment. Reinforcement learning [31] is a type of machine learning algorithms where an agent learns a policy to make decisions by performing actions in an environment to maximize the cumulative reward. In the context of DBMS knobs tuning, RL-based methods [41, 19, 12, 20] treat it as a trial-and-error process based on system feedback and exhibit a degree of adaptability to unknown runtime environments. Although recent studies [2, 9, 14, 24, 36, 42] have enhanced the RL-based methods from various angles, training reinforcement learning models is still significant costly, especially in the field of knobs tuning, where each configuration evaluation for transition collection may require significant time consumption. The offline agent training usually suffers from the poor initial policy since the high-performing configurations are very sparse in the configuration space.
@@ II-B Tuning Hints Extraction
Manuals, web forums and other texts often offer a wealth of domain knowledge about knob settings for DBMSs. DB-BERT [33] leverages BERT to mine these tuning hints in the form of <<knob, rec_value>>, where rec_value is the recommended value for the corresponding knobknob. However, as shown in Table I, since some texts only provide implicit tuning advice such as increasing and decreasing the knob, the extracted tuning hints should contain these tuning direction information. Besides, we should also pay attention to their application condition if we want accurately using the extracted tuning hints. Towards this target, we extend the definition of tuning hint as a recommended adjustment action for a knob under a specific runtime environment (i.e., application condition). Specifically, a condition-aware tuning hint can be described as a triple <<knob, rec_action, condition_vector>>, where condition_vector is a vector describes the applicable runtime environment of rec_action. We will give a detailed description about the definition of condition_vector and how to extract condition-aware tuning hints in Section IV.
@@ II-C LLM-Assisted and LLM-Driven Knobs Tuning
As the pioneering work of NLP-enhanced DBMS tuning methods, DB-BERT [33] first utilizes BERT to extract the tuning hints from texts, and then leverages RL to deal with these tuning hints to generate knob values in an iterative manner. Although DB-BERT can incorporate domain knowledge into the knobs tuning task, however, it totally neglects the application condition of tuning hints, which may mislead the tuning agent. In addition, it is only able to explore a limited range around the recommended knob values by tuning hints, which may miss the opportunities to further improve the performance. Finally, as shown in Figure 1, totally driven by tuning hints may lose competitiveness along with training steps since the tuning hints’ utilities in performance improvement are not static. Considering the advantages of large language models (LLMs) in semantic understanding, reasoning and domain knowledge, several recent works begin to use LLMs for DBMS knobs tuning. We broadly categorize them into LLM-assisted [18, 15] and LLM-driven [11, 21, 32, 13]. The LLM-assisted methods usually employ a LLM to extract or output tuning hints and then conduct a traditional knobs tuning procedure. For instance, GPTuner [18] uses GPT-4 to select performance-critical knobs and extract the suggested value ranges, and then conduct a Bayesian optimization (BO)-based knobs tuning process. However, tuning hints extracted by the LLMs are not always reliable [22], relying solely on GPT’s judgments does not ensure the alignment and accuracy of domain knowledge with the current environment since BO inherently lacks the ability to interact with runtime environment. We also employ LLMs to assist the tuning hints extraction task in DemoTuner. On the other hand, LLM-driven methods design prompts for the whole knobs tuning workflow including knobs selection, knobs range determination and knobs configuration. While promising, relying solely on LLMs to conduct knobs tuning is still full of challenges at the present stage considering the hallucination and domain adaptation issues.
@@ IV-B LLM-based Tuning Hints Extraction
To effectively mine tuning hints described as <<knob, rec_action, condition_vector>> from the collected source texts, we choose to utilize GPT-3.5-Turbo [27] after evaluating a range of advanced large language models (LLMs) considering the tradeoff between costs and practical effectiveness. How to select the best LLM is out of the scope of this paper. Figure 3 illustrates the prompt structure, which includes four critical elements: the model’s role and expertise, the input source text, the output format, and a task description following the chain of thought (CoT) framework [37]. Specifically, the CoT process can be divided into three steps. The first involves the extraction of knobs specific to the target DBMSs such as MySQL or PostgreSQL. The second focuses on extracting recommended values of knobs. In cases where no specific value is provided, the model is required to infer whether the knob’s value should be increased or decreased. The third targets to extract the corresponding circumstances, which can be utilized to identify the application condition vector of this tuning hint. Incorporating the phrase “Let’s think step by step” to split up the task helps guide GPT’s inference process. Besides, placing constraints directly within each step improves GPT’s comprehension and execution by keeping the task and constraints closely aligned. These prompts sharpen GPT’s focus on DBMS knobs tuning and structure the task into clear steps, thus enhancing the extraction accuracy [17]. Notably, GPT’s responses sometimes diverge from the prescribed format and exhibit variability. In such cases, we employ GPT to further analyze the deviant text until it aligns with our specified processing standards.
@@ V-B Discussions About Demonstration Reinforcement Learning
During the training process of DRL, the policy network learns to select actions by optimizing its parameters based on environmental feedback, while the value function estimates the expected rewards for these actions, guiding the policy network through transitions to maximize cumulative rewards. Without domain knowledge, the training process usually has to start from scratch and thus suffers from the cold start problem. Unfortunately, conventional DRL algorithms including Deep Q-Networks (DQN) [25], Policy Gradients [29] and Actor-Critic [23] cannot directly integrate domain knowledge into the training phase. To address this issue, demonstration reinforcement learning algorithms such as DDPGfD (Deep Deterministic Policy Gradient from Demonstrations) have been recently developed [35]. These algorithms can integrate demonstration experiences into the learning process, effectively bridging the gap between pure reinforcement learning and imitation learning.